OpenStudy (anonymous):

Find the derivative of the function. f(x) = 3x2 ln(8x)

OpenStudy (dumbcow):

use product rule and chain rule f(x)=3x, g(h(x)) = 2ln(h), h(x) = 8x chain rule: g(x)' = h(x)'g(h)' product rule: fg(x)' = f'g + fg'

OpenStudy (anonymous):

the problem got typed in wrong it should say 3x^2 ln(8x)

OpenStudy (dumbcow):

ok sorry just change the functions accordingly same steps though

OpenStudy (anonymous):

im still confused about wat u did

OpenStudy (dumbcow):

ok just split it into two parts 3x^2ln(8x) = (3x^2)*(ln(8x)) Now think of it as two different functions multiplied together

OpenStudy (dumbcow):

then apply the product rule call one function f and the other g (f*g)' = f'g +g'f